ansform the given differential equation to find a nontrivial solution such that x(0) = 0. tx + (t - 4)x + x = 0 Click theCan someone help by explaining the correct choice?
Answer #1

ANSWER: 1) sty xk+6*4)=0 ; xft)=cket,cto DETAILS Given differential equation is tall +64-47* ta so tall ttal-43tazo 5 [t2]210) 20 Do [5X)_x.co)-a [s X43] ~4 stte) +)=- 1: 2_25x4) as? xk) –X)sxte) ~45X($)+X) =- | pifferentiating First twotasime Ats